mary has 36 bill in her piers, 10 bills and 20 bills in total 460 dollars,
how many 10 and 20 bills are in her piers

Answer by checkley77(12844)

x+y=36
x=36-y
10x+20y=460
10(36-y)+20y=460
360-10y+20y=460
10y=460-360
10y=100
y=100/10
y=10 twenty dollar bills.
x+10=36
x=36-10
x=26 10 dollar bills.
Proof:
10*26+20*10=460
260+200=460
460=460